A ladder 40 ft long leans against a building and reaches a ledge. The foot of the ladder is 16 ft from the building. How high is the ledge?
Answer by scott8148(6628)

let h=height

by Pythagoras __ h^2+16^2=40^2 __ h^2=40^2-16^2 __ h=8sqrt(21)